Sarah has 7 keys. Two of the keys unlock her desk drawer. If the first key she randomly chooses does not open the lock and she sets it aside, what is the probability that the second key she randomly chooses will unlock the drawer?

1
Expert's answer
2021-05-07T16:42:52-0400

If the first key does not open the lock, then Sarah has 6 keys, of which 2 will open the lock. Then wanted probability is

"p = \frac{2}{6} = \frac{1}{3}"

Answer: "p = \frac{1}{3}"